OpenStudy (anonymous):

I did the long division but how do I explain if 3x+1 is the factor of the dividend ?

OpenStudy (fibonaccichick666):

can you ellaborate a little?

OpenStudy (amistre64):

the remiander will be zero also, if it IS a factor P(x) = (3x+1) Q(x) + R(x) when 3x+1 = 0 ...., x=-1/3 P(-1/3) = (0) Q(-1/3) + R(-1/3) P(-1/3) = R(-1/3) if P(-1/3) = 0, then (3x+1) is a factor

OpenStudy (anonymous):

I posted a picture

OpenStudy (anonymous):

Can you see it ?

OpenStudy (amistre64):

did you attach a file? or a link?

OpenStudy (anonymous):

attached a file i did it form my phone

OpenStudy (amistre64):

no post with pick is present on my end

OpenStudy (fibonaccichick666):

same here

OpenStudy (amistre64):

regardless is the dividend equal to 0 when x=-1/3 ?

OpenStudy (anonymous):

okay let me posted it form the computer one sec i just really need help and i dont get it

OpenStudy (fibonaccichick666):

It doesn't have to equal zero, Though I need to see the whole problem to understand your question

OpenStudy (amistre64):

of course it has to equal zero ... P(x) factors into A(x) Q(x) let A(x) = (3x+1) P(x) = (3x+1) Q(x), when 3x+1 = 0, P(x) = 0

OpenStudy (amistre64):

if P(-1/3) is not equal to 0, then (3x+1) is not a factor

OpenStudy (fibonaccichick666):

not in polynomial division

OpenStudy (amistre64):

im talking of polynomial division ....

OpenStudy (fibonaccichick666):

dividend is just what you are dividing by. Something can be a factor of the dividend and not =0

OpenStudy (fibonaccichick666):

is 3x+1 not a factor of (9x^2-1)?

OpenStudy (fibonaccichick666):

but, I think we need to see the question, I think her question is how to express the remainder

OpenStudy (amistre64):

dividend ------- = quotient divisor

OpenStudy (anonymous):

can you see it ?

OpenStudy (anonymous):

i just attached a file

OpenStudy (amistre64):

(3x+1) is a factor of the dividend, the top ... when 3x+1 = 0

OpenStudy (fibonaccichick666):

ok, so you are trying to answer the second part yea? So the only way a number or polynomial is a factor of another, is if you can divide them and be left with a remainder of zero.

OpenStudy (anonymous):

so i have to divide the answer with 3x+1 ?

OpenStudy (fibonaccichick666):

You cannot need decimals with numbers, and with polynomials, you cannot have any left over

OpenStudy (fibonaccichick666):

no, you are just looking at your answer

OpenStudy (fibonaccichick666):

so you divided by 3x+1 right? Did you have a remainder?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

\[\frac{P(x)}{3x+1}=Q(x)\] \[P(x)=(3x+1)~Q(x)~:~definition ~of~factors\] \[P(-1/3)=(3(-1/3)+1)~Q(-1/3)\] \[P(-1/3)=(-1+1)~Q(-1/3)\] \[P(-1/3)=(0)~Q(-1/3)\] \[P(-1/3)=0\]

OpenStudy (fibonaccichick666):

so, the rule in order for something to be a factor is that there is no remainder ie remainder=0. So, can 3x+1 be a factor if you have remainder 3?

OpenStudy (anonymous):

no it cannot be a factor because i have a remainder of 3

OpenStudy (fibonaccichick666):

Then you have answered your question.

OpenStudy (anonymous):

my answer is 3x+1 is not a factor of the dividend because in order for something to be a factor im supposed to have a remainder of 0 but i have a remainder of 3 that is why 3x+1 is not a factor

OpenStudy (anonymous):

is that correct ? just wanted to double check

OpenStudy (fibonaccichick666):

that's a good answer. Do you understand why you must have a remainder of zero though?

OpenStudy (anonymous):

umm ..i just know its the rule .. thats what you told me lol

OpenStudy (fibonaccichick666):

ok, so let's go with why that is the rule. It plays off of what amistre's saying.

OpenStudy (fibonaccichick666):

sa, if you listed the factors of 8 you'd have, 1,2,4,8 right?

OpenStudy (anonymous):

okay :)

OpenStudy (fibonaccichick666):

if I divide 8 by any of those numbers, will I get a whole number?

OpenStudy (anonymous):

umm no?

OpenStudy (fibonaccichick666):

really? what is 8 divided by 2?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

:D lol

OpenStudy (anonymous):

i just got it yes you get a whole number

OpenStudy (fibonaccichick666):

ok, so ie you have no remainder

OpenStudy (anonymous):

8*2=16

OpenStudy (fibonaccichick666):

8/2=4

OpenStudy (anonymous):

oh i thought you said multiply sorry :S

OpenStudy (fibonaccichick666):

so factor just means that when you divide the big number by a factor, you always get a whole number

OpenStudy (anonymous):

oh okay:)

OpenStudy (fibonaccichick666):

so does it make sense now?

OpenStudy (anonymous):

yeah :D thank you very much !!!!!!